Through (4,2) parrallel to y=-3/4x-5
Found 2 solutions by swincher4391, stanbon:
Answer by swincher4391(1107)

A line that is parallel to this line has the same slope. Thus the slope of a parallel line is -3/4
We have a slope and we have a point. So use point-slope
y -y1 = m(x-x1)
y -2 = (-3/4)(x-4)
y -2 = (-3/4)x + 3
y = (-3/4)x + 5 <---- Answer

Answer by stanbon(75887)

Through (4,2) parallel to y=-3/4x-5
---
Slope must be -3/4 and y = 2 when x = 4
Form y = mx + b
Solve for "b":
2 = (-3/4)4+b
b = 5
----
Equation;
y = (-3/4)x + 5
==================
Cheers,
Stan H.
=========